OpenStudy (anonymous):

A box without cover with a square base sides x and height h has a surface area of 12 m sq. Show that the volume of the box expressed in terms of x can be written as V(x) = x(12-x^2)/4. Determine the biggest volume such a box can have.

OpenStudy (anonymous):

surface area of base=x^2 surface area of sides=hx 4hx+x^2=12, 12-x^2=4hx, (12-x^2)/4x=h volume of box=hx^2 V(x)=x^2((12-x^2)/4x), x(12-x^2)/4

OpenStudy (anonymous):

V(x)=3x-x^3/4 V'(x)=3-3x^2/4 V'(x)=0, 3x^2/4=3, x^2=4, x=2 (-2 is not applicable here) max volume is when x=2, h=1

OpenStudy (anonymous):

Thanks a lot